the sum of two consecutive odd integers is -212. find the two integers
Found 2 solutions by macston, Alan3354:
Answer by macston(5194)

I,I+2 are two consecutive odd integers
I+I+2=-212
2I+2=-212 subtract 2 from each side
2I=-214 divide each side by 2
I=-107 first integer is -107
I=I+2==105 Second integer is -105
CHECK
-107+(-105)=-212
-212=-212


Answer by Alan3354(69443)

the sum of two consecutive odd integers is -212. find the two integers
=================
-212/2 = -106
--> -105 & -107